OpenStudy (anonymous):

Find an equation that models the surface area of a sphere in terms of volume.

OpenStudy (agent0smith):

\[\Large SA = 4 \pi r^2\] \[\Large V = \frac{ 4 }{ 3 }\pi r^3 =\] with a bit of manipulation you can make volume look like this \[\Large V = \frac{ 1 }{ 3 }*(4\pi r^2)*r\]

OpenStudy (anonymous):

Thanks a bunch!!!